# Dualtronics: leveraging both faces of polar semiconductors

### v1

@@ Abstract
Unlike non-polar semiconductors such as silicon, the broken inversion symmetry of the wide bandgap semiconductor gallium nitride leads to a large electronic polarization along a unique crystal axis. This makes the two surfaces of the semiconductor wafer perpendicular to the polar axis dramatically different in their physical and chemical properties. In the last three decades, the cation (gallium) face of gallium nitride has been used for photonic devices such as LEDs and lasers. Though the cation face has also been predominantly used for electronic devices, the anion (nitrogen) face has recently shown promise for high electron mobility transistors due to favorable polarization discontinuities. In this work we show that it is possible to make photonic devices on the cation face, and electronic devices on the anion face, of the same semiconductor wafer. This opens the possibility for dualtronics in polar semiconductors, where electronic, photonic, and acoustic properties can be implemented on opposite faces of the same wafer, dramatically enhancing the functional capabilities of this revolutionary semiconductor family.
@@
Figure 1 shows how dual side epitaxy for dualtronics was achieved. The starting wafer is a high transparency bulk n-type O-doped GaN substrate grown by ammonothermal method [16, 17] with mobile electron concentration $\sim$1018 cm-3, and threading dislocation density $\sim$104 cm-2 with both Ga-polar and N-polar sides chemo-mechanically polished to obtain atomically smooth surfaces. The N-polar surface of the GaN substrate was gallium-bonded to a GaN/sapphire carrier wafer as shown in Figure 1(a). Because the surface bonded to the carrier wafer needs to be used again, we chose gallium bonding to the GaN surface of a GaN/sapphire carrier wafer to ensure desired elements (Ga, N) at this bonding interface. Using molecular beam epitaxy (MBE), we first grew a blue (In,Ga,Al)N LED structure consisting of a quantum well active region inside a pn heterojunction on the Ga-polar face of GaN. The wafer was then detached from the carrier wafer and flipped as shown in Figure 1(b). The Ga was removed from the N-polar side, and the metal-polar p-type GaN surface of the LED side was gallium-bonded on to a clean GaN/sapphire carrier wafer. Subsequently, a N-polar GaN/AlGaN/GaN heterostructure was grown on the N-polar face again by MBE to create a high-electron mobility 2DEG to serve as the conduction channel of the HEMT as shown in Figure 1(c). The chosen growth conditions for the LED and HEMT structures are described in the methods section.
Figure 2: (a) Schematic of the HEMT-LED heterostructures grown on both faces of a single crystal c-plane n-GaN substrate. (b) HAADF-STEM image showing the GaN/Al0.40Ga0.60N/GaN HEMT and (c) atomic resolution image corresponding to the uppermost GaN/Al0.40Ga0.60N heterojunction interface which hosts the 2DEG. (d) iDPC image in the uppermost GaN layer of the HEMT indicating that the nitrogen polarity follows that of the substrate to the surface. (e) HAADF-STEM image showing the LED quantum well, electron blocking layer, cladding and contact layers and (f) atomic resolution image corresponding to the LED In0.07Ga0.93N/In0.17Ga0.83N/In0.07Ga0.93N single quantum well. (g) iDPC image of the p-InGaN contact layers of the LED indicating the metal polarity follows that of the substrate.
@@
This photonic active region was followed by p-type doped layers starting with a 20 nm Al0.09Ga0.91N electron blocking layer with Mg concentration of 3 $\times$ $10^{19}\mathrm{cm}^{-3}$ followed by a 150 nm thick GaN layer with a Mg concentration of 6 $\times$ $10^{18}\mathrm{cm}^{-3}$. This LED heterostructure was capped with p-InGaN contact layers consisting of 40 nm thick In0.02Ga0.98N and 5 nm thick In0.15Ga0.85N doped with Mg at the level of 1 $\times$ $10^{20}\mathrm{cm}^{-3}$ and 5 $\times$ $10^{20}\mathrm{cm}^{-3}$, respectively, to form low resistance ohmic contacts for hole injection into the LED active region, while simultaneously to be transparent to the light emitted from the active region. Figure 2(e) shows a HAADF-STEM image of each of these LED layers indicating a uniform quantum well and the surrounding active region along the plane of the wafer. The atomically resolved image in Figure 2(f) shows the 2.5 nm thick In0.17Ga0.83N quantum well and the InGaN cladding layers. The group III element fractions and doping profiles for both the LED and HEMT structure are presented in the time-of-flight secondary ion mass spectrometry (ToF-SIMS) data shown in Extended Data Fig. 1, which indicate that the measured values are as intended. Importantly, there is no noticeable interdiffusion of dopants or group-III broadening of the metal-polar LED layers, indicating interface and doping control is not affected by the second growth step.
Figure 3: (a)-(d) Device processing flow for the double-sided HEMT-LED. Starting from the as-grown heterostructures, the gray arrows follow the independent processing steps chronologically, with the metal-polar LED being processed after the N-polar HEMT. (e) A three-dimensional representation of the complete device. (f) Optical microscope images of the as-processed sample, focused on the LEDs (left) and the HEMTs (right). The N-polar HEMTs are oriented upwards, forming the uppermost surface. For scale, the diameter of the large LED anode contact is 140 µm. (g) Scanning electron microscope (SEM) images of the HEMTs on the N-polar GaN surface (left) and the LEDs on the metal-polar GaN surface (right).
@@
Figure 4: Device characteristics of the HEMTs and LEDs, operating independently. (a) Drain current (black, solid) and transconductance (gray, solid) of an N-polar HEMT as a function of gate-source voltage, operating at a drain-source voltage of 5 V. The black, dashed line indicates a channel sheet current-density of 1 A/mm. (b) Linear plot showing the family of curves for a HEMT (solid black) for a gate-source voltage ranging from 1.75 (on) to -3.25 V (off). On the same scale, the linear current-voltage characteristics of a 400 µm diameter LED is shown as well (blue, solid). The dimensions of the measured HEMTs are LSD = 4 µm, LG = 1.5 µm, and WG = 50 µm. (c) Semi-log plots showing the drain current (black, solid) and gate current (black, dashed) versus gate-source voltage, for a drain-source voltage of 5 V. Here the transistor current corresponds to the left vertical axis. The horizontal, black, dashed line indicates a channel sheet current-density of 1 A/mm. Similarly, corresponding to the right vertical axis is the LED current (blue, solid) as a function of forward bias for a 400 µm diameter device. (d) Electroluminescence spectra of a metal-polar, 400 µm diameter LED. The injection current density ranges from 1 to 100 A/cm2. The top-right inset shows a camera image of the sample with an LED in the on-state.
@@
Figure 5: (a) Energy-band diagrams of the HEMT and LED indicating the off and on-states. (b) Circuit schematic of the monolithic HEMT-LED, taking into account the back-gating effect of the conductive GaN substrate. (c)-(d) Monolithic switching measurements, modulating between on-off (c) and between bright and dim modes (d). The gate-source modulation voltage is shown in red, the photodiode voltage while modulating the LED in solid blue, and the background (LED off state) photodiode voltage in dashed blue.
@@
Figure 3(d) shows the completed dualtronic structure after removal of the protective photoresist coating, and Figure 3(e) shows a cross-section of it flipped upside down. This schematic emphasizes the single-crystal nature of the whole structure, across the bulk of the wafer indicated by the wavy break in the thick GaN layer between the HEMT on the top and the LED on the bottom. Figure 3(f) shows an optical microscope image focused on the top HEMT surface on the left, and the LED focused on the right. The energy bandgap of the bulk wafer between the HEMT and LED mesas is $E_{g1}=3.4$ eV of GaN, which is optically transparent to visible light, allowing for imaging of the LEDs through the wafer from the HEMT side as seen in Figure 3(f). This also emphasizes that the transistor layers are transparent to visible wavelengths, and are therefore attractive for dualtronic integration with LEDs as transparent thin-film transistors (TFTs) but, as seen later, with dramatically higher performance than existing TFTs made of oxide or organic semiconductors. Figure 3(g) shows the scanning electron microscope (SEM) images of the HEMTs on the N-polar GaN surface (left) and the LEDs on the metal-polar GaN surface (right).
@@
Figure 4(a) shows the measured transfer characteristic and transconductance of the N-polar HEMT. A threshold voltage of $V_{T}\approx-3$ V is measured with an on-current drive exceeding 1 A/mm at a drain voltage of $V_{DS}=5$ V and gate voltage $V_{GS}\approx 3$ V. The transistor transconductance $g_{m}=\partial I_{D}/\partial V_{GS}$ peaks at $V_{GS}\approx-1.6$ V. Figure 4(b) shows the measured $I_{D}-V_{DS}$ output characteristics of the HEMT for various gate voltages overlaid with the LED current-voltage characteristics. The transistor delivers enough output current to drive the LED with variable drive currents controlled by the gate voltage up to 50 mA or more. Figure 4(c) shows the LED current and the HEMT current for $V_{DS}=5$ V in the logarithmic scale indicating that the HEMT can control the LED current and hence its optical output intensity over several orders of magnitude. Figure 4(d) shows the measured output spectra of the LED for injection current densities ranging from 1 to 100 A/cm2. The emission intensity increases with the injection current density, and the peak emission wavelength blue-shifts from $\lambda\approx 470$ nm at 1 A/cm2 to $\lambda\approx 450$ nm at 100 A/cm2 in the blue regime of the visible spectrum. The blueshift is caused by screening of the internal polarization fields which result in a quantum confined Stark effect (QCSE), as well as the Burstein-Moss effect. The inset in Figure 4(d) shows the bright blue electroluminescent glow of the blue LED slightly below the center of the wafer on the metal-polar side. The output and transfer characteristics of the HEMTs before and after the fabrication of the LEDs is shown in the Extended Data Fig. 5. It was confirmed that after the LED processing, there was minimal degradation of the HEMT with a less then 0.3 V threshold voltage shift and negligible change in the output and gate leakage current. This minimal degradation highlights the feasibility of using double-side epitaxy and processing for reliable heterogeneous device integration on both crystal faces.
@@
Figure 5(c) shows that the photovoltage signal follows the gate voltage as it is swept between $-3.0\leq V_{\mathrm{GS}}\leq 3.0$ V with the drain voltage set at a voltage of 5.9 V and the backgate voltage set at a voltage of 1 V, over millisecond timescale across its threshold voltage from on state to off state. Figure 5(d) shows that when the gate voltage of the HEMT is modulated between $2.0\leq V_{\mathrm{GS}}\leq 3.0$ V when it is in its on state with $V_{\mathrm{GS}}>V_{\mathrm{T}}$, the light emitted by the LED follows this gate modulation, translating the electronic signal in the gate of the HEMT into a photonic bright and dim signal output from the LED on the other side of the wafer.
@@
The observations described here thus prove that the concept of dualtronics is feasible, opening the gates for many interesting possibilities. This work realized an electronics fabric on the N-polar face and a photonic fabric on the metal-polar face of the same wafer. This dualtronic combination is of immediate interest for combining microLEDs on the photonic side with transparent TFTs on the same GaN wafer on the other side. This monolithic convergence of devices reduces the number of components required for microLEDs with the potential for dramatic area and cost savings due to efficient use of the substrate real-estate. Moreover, monolithic integrating schemes combining HEMTs and LEDs on a single substrate face rely on multiple epitaxial layers that need to be either selectively removed and/or regrown in order to expose the LED heterostructure buried underneath the HEMT structure or vise versa [23, 24, 25, 26, 27]. Exposing the buried heterostructure by dry etching induces plasma damage, leading to diminished light emission and degradation of contacts. Similarly, with a selective growth method, the ex-situ nature of regrowth results in a poor growth interface, increasing leakage pathways. These issues are all avoided with the dualtronics integration scheme.
@@
But the dualtronics concept extends to several exciting new opportunities. Laser diodes, semiconductor optical amplifiers (SOAs), and electro-optical modulators can be realized on one side of the wafer, and transistors on the other side, dramatically decreasing the components and chips desired in photonic integrated circuits. The combination of photodetectors and LEDs or Lasers on both sides can enable an on-chip photonic communication and computation system. RF transistor power amplifiers (PAs) for transmit part of communication systems can be realized on one polarity, and low-noise amplifiers (LNAs) for the receive end of communication systems can enable integrated transceivers in a combined transmit/receive module of smaller form factor than existing systems. The combination of n-channel transistors on one polarity with p-channel transistors on the opposite polarity can enable new forms of complementary transistor circuit topologies connected by through-vias through the substrate. Such dualtronic devices can take advantage of the wide bandgap nature of the polar nitride semiconductors for new forms of power electronics and RF electronics. These, and several allied possibilities can allow for creation and manipulation of electrons and photons on the opposite faces of the same wafer to achieve new functionalities.
@@
The ultra wide bandgap polar semiconductor AlN boasts large electro-acoustic coupling, which makes it the preferred material today for acoustic wave RF filters. Dualtronics can thus take advantage of this property of the polar nitride semiconductors to combine sonar (by sound waves), radar (microwaves), and lidar (light) on the same platform. The efficient use of the substrate surfaces eliminates wasted space, reduces the energy and material costs of producing multiple wafers, and thus should be of high interest for future technologies well beyond the particular semiconductor materials discussed here.

### v2

# Leveraging both faces of polar semiconductor wafers for functional devices

@@ Abstract
Unlike non-polar semiconductors such as silicon, the broken inversion symmetry of the wide bandgap semiconductor gallium nitride leads to a large electronic polarization along a unique crystal axis[1]. This makes the two surfaces of the semiconductor wafer perpendicular to the polar axis dramatically different in their physical and chemical properties[2]. In the last three decades, the cation (gallium) face of gallium nitride has been used for photonic devices such as LEDs and lasers[3, 4, 5]. Though the cation face has also been predominantly used for electronic devices, the anion (nitrogen) face has recently shown promise for high electron mobility transistors due to favorable polarization discontinuities[6]. In this work we introduce dualtronics, showing that it is possible to make photonic devices on the cation face, and electronic devices on the anion face, of the same semiconductor wafer. This opens the possibility for leveraging both faces of polar semiconductors in a single structure, where electronic, photonic, and acoustic properties can be implemented on opposite faces of the same wafer, dramatically enhancing the functional capabilities of this revolutionary semiconductor family.
@@
Figure 1 shows how dual side epitaxy for dualtronics was achieved. The starting wafer is a high transparency bulk n-type O-doped GaN substrate grown by ammonothermal method [22, 23] with mobile electron concentration $\sim$1018 cm-3, and threading dislocation density $\sim$104 cm-2 with both the Ga-polar and N-polar sides chemo-mechanically polished to obtain atomically smooth surfaces. The N-polar surface of the GaN substrate was gallium-bonded to a GaN/sapphire carrier wafer as shown in Figure 1(a). Because the surface bonded to the carrier wafer needs to be used again, we chose gallium bonding to the GaN surface of a GaN/sapphire carrier wafer to ensure desired elements (Ga, N) at this bonding interface. Using molecular beam epitaxy (MBE), we first grew a blue (In,Ga,Al)N LED structure consisting of a quantum well active region inside a pn heterojunction on the Ga-polar face of GaN. The wafer was then detached from the carrier wafer and flipped as shown in Figure 1(b). The Ga was removed from the N-polar side, and the metal-polar p-type GaN surface of the LED side was gallium-bonded on to a clean GaN/sapphire carrier wafer. Subsequently, a N-polar GaN/AlGaN/GaN heterostructure was grown on the N-polar face again by MBE to create a high-electron mobility 2DEG to serve as the conduction channel of the HEMT as shown in Figure 1(c). The chosen growth conditions for the LED and HEMT structures are described in the methods section.
Figure 2: STEM imaging of dualtronic epitaxial heterostructure. (a) Schematic of the HEMT-LED heterostructures grown on both faces of a single crystal c-plane n-GaN substrate. (b) HAADF-STEM image showing the GaN/Al0.40Ga0.60N/GaN HEMT and (c) atomic resolution image corresponding to the uppermost GaN/Al0.40Ga0.60N heterojunction interface which hosts the 2DEG. (d) iDPC image in the uppermost GaN layer of the HEMT indicating that the nitrogen polarity follows that of the substrate to the surface. (e) HAADF-STEM image showing the LED quantum well, electron blocking layer, cladding and contact layers and (f) atomic resolution image corresponding to the LED In0.07Ga0.93N/In0.17Ga0.83N/In0.07Ga0.93N single quantum well. (g) iDPC image of the p-InGaN contact layers of the LED indicating the metal polarity follows that of the substrate.
@@
This photonic active region was followed by p-type doped layers starting with a 20 nm Al0.09Ga0.91N electron blocking layer with a Mg concentration of 3 $\times$ $10^{19}\mathrm{cm}^{-3}$ followed by a 150 nm thick GaN layer with a Mg concentration of 6 $\times$ $10^{18}\mathrm{cm}^{-3}$. This LED heterostructure was capped with p-InGaN contact layers consisting of 40 nm thick In0.02Ga0.98N and 5 nm thick In0.15Ga0.85N doped with Mg at the level of 1 $\times$ $10^{20}\mathrm{cm}^{-3}$ and 5 $\times$ $10^{20}\mathrm{cm}^{-3}$, respectively, to form low resistance ohmic contacts for hole injection into the LED active region, while simultaneously to be transparent to the light emitted from the active region. Figure 2(e) shows a HAADF-STEM image of each of these LED layers indicating a uniform quantum well and the surrounding active region along the plane of the wafer. The atomically resolved image in Figure 2(f) shows the 2.5 nm thick In0.17Ga0.83N quantum well and the InGaN cladding layers. The group III element fractions and doping profiles for both the LED and HEMT structure are presented in the time-of-flight secondary ion mass spectrometry (ToF-SIMS) data shown in Extended Data Fig. 2, which indicate that the measured values are as intended. Importantly, there is no noticeable interdiffusion of dopants or group-III broadening of the metal-polar LED layers, indicating interface and doping control is not affected by the second growth step.
Figure 3: Fabrication and imaging of dualtronic device. (a)-(d) Device processing flow for the double-sided HEMT-LED. Starting from the as-grown heterostructures, the gray arrows follow the independent processing steps chronologically, with the metal-polar LED being processed after the N-polar HEMT. (e) A three-dimensional representation of the complete device. (f) Optical microscope images of the as-processed sample, focused on the LEDs (right) and the HEMTs (left). The N-polar HEMTs are oriented upwards, forming the uppermost surface. For scale, the diameter of the large LED anode contact is 140 µm. (g) Scanning electron microscope (SEM) images of the HEMTs on the N-polar GaN surface (bottom) and the LEDs on the metal-polar GaN surface (top).
@@
Figure 4: Device characteristics of the HEMTs and LEDs, operating independently. (a) Normalized drain current (black, solid) and transconductance (gray, solid) of an N-polar HEMT as a function of gate-source voltage, operating at a drain-source voltage of 5 V. (b) Linear plot showing the family of curves for a HEMT (solid black) for a gate-source voltage ranging from 1.75 (on) to -3.25 V (off). On the right axis, the linear current-voltage characteristics of a 400 µm diameter LED (blue, solid) and unnormalized output characteristics of a HEMT are shown as well. The dimensions of the measured HEMTs are LSD = 4 µm, LG = 1.5 µm, and WG = 50 µm. (c) Semi-log plots showing the unnormalized drain current (black, solid) and gate current (black, dashed) versus gate-source voltage, for a drain-source voltage of 5 V. Here the transistor current corresponds to the left vertical axis. The horizontal, black, dashed line indicates a normalized channel sheet current-density of 1 A/mm. Similarly, corresponding to the right vertical axis is the LED current (blue, solid) as a function of forward bias for a 400 µm diameter device. (d) Electroluminescence spectra of a metal-polar, 400 µm diameter LED. The injection current density ranges from 1 to 140 A/cm2. The top-right inset shows a camera image of the sample with an LED in the on-state.
@@
Figure 5: Monolithic HEMT-LED switching measurements. (a) Energy-band diagrams of the HEMT and LED indicating the off and on-states. (b) Circuit schematic of the monolithic HEMT-LED, taking into account the back-gating effect of the conductive GaN substrate. (c)-(d) Monolithic switching measurements, modulating between on-off (c) and between bright and dim modes (d). The gate-source modulation voltage is shown in red, the photodiode voltage while modulating the LED in solid blue, and the background (LED off state) photodiode voltage in dashed blue.
@@
Figure 3(d) shows the completed dualtronic structure after removal of the protective photoresist coating, and Figure 3(e) shows a cross-section of it flipped upside down. This schematic emphasizes the single-crystal nature of the whole structure, across the bulk of the wafer indicated by the wavy break in the thick GaN layer between the HEMT on the top and the LED on the bottom. Figure 3(f) shows an optical microscope image focused on the top HEMT surface on the left, and the LED focused on the right. The energy bandgap of the bulk wafer between the HEMT and LED mesas is $E_{g1}=3.4$ eV of GaN, which is optically transparent to visible light, allowing for imaging of the LEDs through the wafer from the HEMT side as seen in Figure 3(f). This also emphasizes that the transistor layers are transparent to visible wavelengths, and are therefore attractive for dualtronic integration with LEDs as transparent thin-film transistors (TFTs) but, as seen later, with dramatically higher performance than existing TFTs made of oxide or organic semiconductors. Figure 3(g) shows the scanning electron microscope (SEM) images of the HEMTs on the N-polar GaN surface (bottom) and the LEDs on the metal-polar GaN surface (top).
@@
Figure 4(a) shows the measured transfer characteristic and transconductance of the N-polar HEMT. A threshold voltage of $V_{T}\approx-3$ V is measured with an on-current drive exceeding 1 A/mm at a drain voltage of $V_{DS}=5$ V and gate voltage $V_{GS}\approx 3$ V. The transistor transconductance $g_{m}=\partial I_{D}/\partial V_{GS}$ peaks at $V_{GS}\approx-1.6$ V. Figure 4(b) shows the measured $I_{D}-V_{DS}$ output characteristics of the HEMT for various gate voltages overlaid with the LED current-voltage characteristics. The transistor delivers enough output current to drive the LED with variable drive currents controlled by the gate voltage up to 50 mA or more. Figure 4(c) shows the LED current and the HEMT current for $V_{DS}=5$ V in the logarithmic scale indicating that the HEMT can control the LED current and hence its optical output intensity over several orders of magnitude. Figure 4(d) shows the measured output spectra of the LED for injection current densities ranging from 1 to 140 A/cm2. The emission intensity increases with the injection current density, and the peak emission wavelength blue-shifts from $\lambda\approx 470$ nm at 1 A/cm2 to $\lambda\approx 450$ nm at 140 A/cm2 in the blue regime of the visible spectrum. The blueshift is caused by screening of the internal polarization fields which result in a reduced quantum confined Stark effect (QCSE), as well as the Burstein-Moss effect. The inset in Figure 4(d) shows the bright blue electroluminescent glow of the blue LED slightly below the center of the wafer on the metal-polar side. The output and transfer characteristics of the HEMTs before and after the fabrication of the LEDs is shown in the Extended Data Fig. 6. It was confirmed that after the LED processing, there was minimal degradation of the HEMT with a less then 0.3 V threshold voltage shift and negligible change in the output and gate leakage current. This minimal degradation highlights the feasibility of using double-side epitaxy and processing for reliable heterogeneous device integration on both crystal faces.
@@
Figure 5(c) shows that the photovoltage signal follows the gate voltage as it is swept between $-3.0\leq V_{\mathrm{GS}}\leq$ 3.6 V with the drain voltage set at a voltage of 5.9 V and the backgate voltage set at a voltage of 1 V, over millisecond timescale across its threshold voltage from on state to off state. Figure 5(d) shows that when the gate voltage of the HEMT is modulated between $2.0\leq V_{\mathrm{GS}}\leq 3.0$ V when it is in its on state with $V_{\mathrm{GS}}>V_{\mathrm{T}}$, the light emitted by the LED follows this gate modulation, translating the electronic signal in the gate of the HEMT into a photonic bright and dim signal output from the LED on the other side of the wafer. The on-off modulation of the dualtronic device can be improved by utilizing similar concepts as used for traditional LEDs. [29, 30]
@@
The observations described here thus prove that the concept of dualtronics is feasible, opening the gates for many interesting possibilities. This work realized an electronics fabric on the N-polar face and a photonic fabric on the metal-polar face of the same wafer. This dualtronic combination is of immediate interest for combining microLEDs on the photonic side with transparent TFTs on the same GaN wafer on the other side. This monolithic convergence of devices reduces the number of components required for microLEDs with the potential for dramatic area and cost savings due to efficient use of the substrate real-estate. Moreover, monolithic integrating schemes combining transistors and LEDs on a single substrate face rely on multiple epitaxial layers that need to be either selectively removed and/or regrown in order to expose the LED heterostructure buried underneath the transistor structure or vise versa [31, 32, 33, 34, 35, 36, 37, 38]. Exposing the buried heterostructure by dry etching induces plasma damage, leading to diminished light emission and degradation of contacts. Similarly, with a selective growth method, the ex-situ nature of regrowth results in a poor growth interface, increasing leakage pathways. These issues are all avoided with the dualtronics integration scheme.
@@
But the dualtronics concept extends to several exciting new opportunities. The metal-polar face of the substrate, to take advantage of high emission efficiencies, can be utilized for any optoelectronic device like laser diodes, semiconductor optical amplifiers (SOAs), and electro-optical modulators, while transistors or photodetectors are fabricated on the N-polar face. This full utilization of the substrate dramatically decreases the number of components and chips needed in photonic integrated circuits. For other applications, both GaN polarities can be used. For instance, RF transistor power amplifiers (PAs) for the transmit part of communication systems can be realized on one polarity, and low-noise amplifiers (LNAs) for the receive end of communication systems can enable integrated transceivers in a combined transmit/receive module of smaller form factor than existing systems. The combination of n-channel transistors on one polarity with p-channel transistors on the opposite polarity can enable new forms of complementary transistor circuit topologies connected by through-vias through the substrate. Such dualtronic devices can take advantage of the wide bandgap nature of the polar nitride semiconductors for new forms of power electronics and RF electronics. These, and several allied possibilities, can allow for the creation and manipulation of electrons and photons on the opposite faces of the same wafer to achieve new functionalities.
@@
The ultra wide bandgap polar semiconductor AlN boasts large electro-acoustic coupling, which makes it the preferred material today for acoustic wave RF filters. Dualtronics can thus take advantage of this property of the polar nitride semiconductors to combine sonar (by sound waves), radar (microwaves), and lidar (light) on the same platform.
The efficient use of the substrate surfaces eliminates wasted space, reduces the energy and material costs of producing multiple wafers, and thus should be of high interest for future technologies well beyond the particular polar semiconductor materials discussed here.